# Rule-driven News Captioning

### 11 Mar 2024 v2

@@ IV-C Ablative Analysis
P4(10 $\sim$ 12) 9.38 13.57 24.32 72.76
Effectiveness of Embedding Named Entities (Q2). During the rule construction, we employ named entities extracted from news articles to replace generic objects in Algorithm 1. In order to verify its effectiveness, we design the variant Non-Entity, which removes the process of named entity extraction and replacement (Eq. 5) from our full model. Namely, we directly use $F^{\prime}$ in Eq. 4 as the rule to guide the model learning. We encode $F^{\prime}$ using CLIP and transform it into the rule vector $f^{\prime}$ through a linear layer. We rewrite Eq. 7 as:
$f^{\prime}=\mathrm{Linear}(\mathrm{CLIP_{text}}(F^{\prime}))$ (15)
@@
We observe that the deep layers P4 can achieve the better performances than the shallow layers P1 or P2, across all metrics on two datasets. Especially, in the CIDEr score, our method (P4) can achieve 72.92% on GoodNews and 72.76% on NYTimes800k, with a relative improvement of 6.16% and 2.75% over P1. It illustrates that the deeper encoder layers are more suitable for embedding news-aware semantic rules in image caption generation.
@@ IV-D Qualitative Analysis
Figure 5: Qualitative results of our method on the GoodNews dataset. We provide the constructed news-aware semantic rule in each example. Non-Rule refers to the caption generated without using news-aware semantic rule. Tell uses the byte-pair-encoding transformer to generate captions [3]. GT is the ground-truth caption. Correct sentences are marked in green, while incorrect ones are marked in red.
@@
    Our model can capture core events in images to construct the news-aware semantic rule, which is used to accurately depict event-related information to produce captions. For example, in case (a), Both Non-Rule and Tell cannot effectively capture the main action “distributing” in the image, so their generated captions are insufficient to accurately describe the image content. In contrast, our method effectively extracts the key action “distributing” and its associated entities using a news-aware semantic rule. We then incorporate this rule into BART, resulting in the generation of accurate news captions.
@@
    During the construction of semantic rules, each role can be correlated with corresponding named entities to facilitate the news caption generation. In case (c), our approach connects the role of “Agent” and the named entity “Jon Kyl” by providing explicit affiliations for generic objects and named entities, thereby generating the correct description, “‘Most of these … should be eliminated,’ said Jon Kyl, a retired Republican senator from Arizona”. In contrast, due to the lack of rule guidance, both Non-Rule and Tell generate the incorrect named entities “Ron Wyden” and “Harry Reid”, respectively.
@@
    We notice a failure case in case (d), where our model failed to output information about “Residents of West Pittston”. The reason is that our semantic rule is designed to capture key events from image, so our model only outputs information “The Susquehanna River overflowed its banks in West Pittston”. In future work, integrating multi-modal knowledge from both images and news articles into rules may have a positive impact on improving the performance of our model.

### 14 Mar 2024 v3

@@ IV-C Ablative Analysis
P4(10 $\sim$ 12) 9.41 13.10 24.42 72.29
Effectiveness of Embedding Named Entities (Q2). During the rule construction, we employ named entities extracted from news articles to replace generic objects in Algorithm 1. In order to verify its effectiveness, we design the variant Non-Entity, which removes the process of named entity extraction and replacement (Eq. 5) from our full model. Namely, we directly use $F^{\prime}$ in Eq. 4 as the rule to guide the model learning. We encode $F^{\prime}$ using CLIP and transform it into the rule vector $z^{f^{\prime}}$ through a linear layer. We rewrite Eq. 7 as:
$z^{f^{\prime}}=\mathrm{Linear}(\mathrm{CLIP_{text}}(F^{\prime}))$ (15)
@@
We observe that the deep layers P4 can achieve the better performances than the shallow layers P1 or P2, across all metrics on two datasets. Especially, in the CIDEr score, our method (P4) can achieve 71.58% on GoodNews and 72.29% on NYTimes800k, with a relative improvement of 4.82% and 2.28% over P1. It illustrates that the deeper encoder layers are more suitable for embedding news-aware semantic rules in image caption generation.
@@ IV-D Qualitative Analysis
Figure 5: Qualitative results of our method. We provide the constructed news-aware semantic rule in each example. Non-Rule refers to the caption generated without using news-aware semantic rule. Tell uses the byte-pair-encoding transformer to generate captions [3]. GT is the ground-truth caption. Correct sentences are marked in green, while incorrect ones are marked in red.
@@
    During the construction of semantic rules, we link roles with named entities to guide the caption generation. In case (a), our approach links the role of “Agent” and the named entity “Lars Bender” by providing explicit affiliations for generic objects and named entities, thereby generating the correct description, “Lars Bender scored the winning goal for Germany in the 80th minute.” In contrast, due to the lack of rule guidance, both Non-Rule and Tell generate the incorrect named entities “Lukas Podolski”.
@@
    Our model can capture core events in images to construct the news-aware semantic rule, which is used to accurately depict event-related information to produce captions. For example, in case (c), Non-Rule cannot effectively capture the main action “performing” in the image, while Tell cannot correctly recognize the named entity “Femi Kuti” associated with the action. So their generated captions are insufficient to accurately describe the image content. In contrast, our method effectively extracts the key action “performing” and its associated entities “Femi Kuti” using a news-aware semantic rule. We then incorporate this rule into BART, resulting in the generation of accurate news captions.
@@
    We notice a failure case in case (d), where our model failed to output information about “Catholic Charities of the Diocese”. The reason is that our semantic rule is designed to capture key events from image, so our model only outputs information “A fire destroyed the largest halfway house in Paterson, N.J., on Saturday.” In future work, integrating multi-modal knowledge from both images and news articles into rules may have a positive impact on improving the performance of our model.